OpenStudy (anonymous):

I don't understand this please help... A collection of dimes and nickels is worth $6.90. There are 96 coins in all. How many of each coin are there? dimes nickels

OpenStudy (danjs):

welcome to open study

OpenStudy (anonymous):

Hai

OpenStudy (danjs):

You need to set up 2 equations here

OpenStudy (danjs):

Total Coins = Dimes + Nickles Let D = # Dimes and N = # Nickles 96 = D + N

OpenStudy (anonymous):

0.10x + 0.05y=6.90 x+y=96

OpenStudy (danjs):

right those are the ones

OpenStudy (anonymous):

but what do I do now ?

OpenStudy (danjs):

Need to find the value for X and Y that will make both of those equations true, there are a couple different ways to go about that

OpenStudy (danjs):

We will do a substitution here....Solve one equation for X first X + Y = 96 so X = 96 - Y

OpenStudy (danjs):

good?

OpenStudy (anonymous):

Mhm

OpenStudy (danjs):

Now, replace X in the other equation by X = 96 - Y 0.10x + 0.05y = 6.90 0.10*(96-y) + 0.05y = 6.90

OpenStudy (danjs):

that make sense there?

OpenStudy (anonymous):

That would make 9.60-0.10y+0.05=6.90

OpenStudy (danjs):

yes, just forgot the y after 0.05

OpenStudy (anonymous):

Oh yeah, sorry

OpenStudy (anonymous):

Combine like terms ?

OpenStudy (danjs):

9.60- 0.10y+ 0.05y = 6.90 add together the y's 9.60 - 0.05y = 6.90 subtract 9.60 -0.05y = -2.70 Divide by -0.05 y = 54

OpenStudy (danjs):

y here is the number of nickles # nickles = y = 54

OpenStudy (anonymous):

nickles=54 and dimes=42? Correct ?

OpenStudy (danjs):

You can test it, Total coins = 96 = 54 + 42 ? 6.90 = 0.05*54 + 0.10*42 ?? if both are true, then you have the right answer

OpenStudy (anonymous):

Thanks A lot =D

OpenStudy (danjs):

welcome, anytime